# Versor: A Geometric Sequence Architecture Enhanced Scale Generalization and Interpretability via Conformal Algebra

### v1

@@ F.1.1 Quantitative Ablation Study
Analysis: As shown in Table 4, Manifold Normalization is critical for convergence; removing it leads to immediate divergence due to scale explosion. Interestingly, removing the Recursive Rotor structure (replacing it with a standard RNN update on the manifold) yields stable but less interpretable results. The Full Versor outperforms the Standard Transformer baseline significantly in terms of Mean MSE (3.44 vs 6.38).
@@ F.1.4 Density Stress Test Details
Versor (Base) 16 Lanes 0.20M 5.210
Versor (Multi-Channel) 16 Lanes 0.85M 3.067
@@
The Multi-Channel Versor scales with internal lane capacity. Note: The Multi-Channel variant tested here uses cross-channel mixing (allowing geometric interactions between channels), which increases parameters to 0.85M compared to the block-diagonal Base model (0.20M). This is distinct from the efficiency comparison in Section 6.3, where the block-diagonal Multi-Channel architecture is compared against a fully-dense Standard baseline. Here, we test whether adding geometric mixing between channels improves capacity. By matching the baseline’s 16-lane architecture and utilizing Clifford-equivariant mixing (Appendix D.1), the model achieves a 41% error reduction (Table above), confirming that geometric channel interactions provide representational benefits beyond mere parameter depth.
@@ F.2 Detailed Multimodal Results
In Section 6.4, we summarized the performance of Versor across NLP, Vision, and Graph tasks. Here, we provide the full aggregated statistics across 5 random seeds (42–46) to evaluate stability and convergence.
@@ G.2.1 Robotics: The “Loop Closure” Problem
In SLAM (Simultaneous Localization and Mapping), a robot must deduce its location from a sequence of movements. Standard RNNs suffer from “drift” where accumulated errors lead to invalid rotation matrices (shearing). Versor’s manifold constraint ensures that the estimated pose is always a valid element of $SE(3)$. The Recursive Rotor Accumulator is isomorphic to the standard Lie Group Integrators used in control theory, but learnable. This could enable end-to-end learning of robust odometry from noisy IMU data.
@@ G.2.2 Biochemistry: Protein Folding beyond AlphaFold
Proteins are chains of amino acids where the relative orientation of peptide planes determines the structure. AlphaFold [15] uses “Invariant Point Attention” (IPA) to explicitly model these frames. Versor offers a more natural representation: modeling the protein backbone effectively as a “snake” of spinors. Our bivector attention mechanism (GPA) naturally detects “contacts” (scalar proximity) and “relative alignment” (bivector torque) between residues, potentially simplifying the AlphaFold architecture by replacing complex frame-update modules with native Clifford layers.
@@ G.2.3 Relativistic Physics
The algebra $\mathcal{C}l_{4,1}$ used in this paper is the Space-Time Algebra (STA) generally used in special relativity. Unlike standard Euclidean networks, Versor can naturally model Lorentz boosts (rotations involving time $e_{t}$). This suggests applications in High Energy Physics (HEP), such as jet tagging at the Large Hadron Collider, where particle decay products must be analyzed in a frame-independent manner; or modelling of geometries with Lorentzian signatures as suggested in developments of [14].
@@ G.3 Comparison with Quaternion RNNs (QRNN)
A valid question one may have is the relationship between Versor and Quaternion Neural Networks (QRNNs).
@@
Relationship to Dual Quaternions: In robotics, Dual Quaternions ($\mathbb{DH}$) are often used to model rigid body motions (screws). It is important to note that Dual Quaternions are also a subalgebra of CGA. Specifically, the even subalgebra of $\mathcal{C}l_{4,1}$ contains elements isomorphic to dual quaternions. However, existing “Dual Quaternion RNNs” often implement the algebra via $4\times 4$ matrix representations or dual-number libraries which are computationally heavy. Versor’s bit-masked formulation offers a more efficient execution model for this same algebraic structure.

### v2

@@ F.1.1 Quantitative Ablation Study
Result Analysis. As shown in Table 7, Manifold Normalization is critical for convergence; removing it leads to immediate divergence due to scale explosion. Interestingly, removing the Recursive Rotor structure (replacing it with a standard RNN update on the manifold) yields stable but less interpretable results. The Full Versor outperforms the Standard Transformer baseline significantly (3.44 vs 6.38 Mean MSE).
@@ F.1.4 Density Stress Test Details
Versor (Base) 1 Lane 0.007M 5.210
Versor (Multi-Channel) 16 Lanes 1.1M 3.067
@@
The Multi-Channel Versor scales with internal lane capacity. Note: The Multi-Channel variant tested here uses cross-channel mixing (allowing geometric interactions between channels), which increases parameters to 0.85M compared to the block-diagonal Base model (0.20M). This is distinct from the efficiency comparison in Section 6.3, where the block-diagonal Multi-Channel architecture is compared against a fully-dense Standard baseline. Here, it is tested whether adding geometric mixing between channels improves capacity. By matching the baseline’s 16-lane architecture and utilizing Clifford-equivariant mixing (Appendix D.1), the model achieves a 41% error reduction (Table above), confirming that geometric channel interactions provide representational benefits beyond mere parameter depth.
@@ F.2 Detailed Multimodal Results
In Section 6.4, the performance of Versor across NLP, Vision, and Graph tasks was summarized. Here, the full aggregated statistics across 5 random seeds (42–46) are provided to evaluate stability and convergence.
@@ G.2.1 Robotics: The “Loop Closure” Problem
In SLAM (Simultaneous Localization and Mapping), a robot must deduce its location from a sequence of movements. Standard RNNs suffer from “drift” where accumulated errors lead to invalid rotation matrices (shearing). Versor’s manifold constraint ensures that the estimated pose is always a valid element of $SE(3)$. In the verified odometry benchmark, the Recursive Rotor Accumulator (RRA) reduced estimation error by 22% compared to a GRU baseline (MSE $0.000195$ vs. $0.000249$) and achieved a 64% reduction in manifold drift ($0.0044$ vs. $0.0124$). This results from RRA being isomorphic to standard Lie Group Integrators used in control theory, enabling end-to-end learning of robust odometry from noisy IMU sensor data without the “Euclidean Bottleneck” of shearing rotations.
Figure 5: Robotics Odometry Benchmark: Versor (RRA) vs. GRU Baseline. Left: Training loss convergence. Right: Manifold drift comparison, showing Versor’s native ability to maintain $SE(3)$ validity.
@@ G.2.2 Biochemistry: Protein Folding beyond AlphaFold
Proteins are chains of amino acids where the relative orientation of peptide planes determines the structure. AlphaFold [16] uses “Invariant Point Attention” (IPA) to explicitly model these frames. Versor offers a more natural representation: modeling the protein backbone effectively as a “snake” of spinors. The bivector attention mechanism (GPA) naturally detects “contacts” (scalar proximity) and “relative alignment” (bivector torque) between residues, potentially simplifying the AlphaFold architecture by replacing complex frame-update modules with native Clifford layers.
@@ G.2.3 Relativistic Physics
The algebra $\mathcal{C}l_{4,1}$ used in this paper is the Space-Time Algebra (STA) generally used in special relativity. Unlike standard Euclidean networks, Versor can naturally model Lorentz boosts (rotations involving time $e_{t}$). This suggests applications in High Energy Physics (HEP), such as jet tagging at the Large Hadron Collider, where particle decay products must be analyzed in a frame-independent manner; or modeling of geometries with Lorentzian signatures as suggested in developments of [14].
@@ G.3 Comparison with Quaternion RNNs (QRNN)
The relationship between Versor and Quaternion Neural Networks (QRNNs) warrants investigation.
@@
Applications Attitude Control, Color Images Dynamics, Robotics, Physics
Relationship to Dual Quaternions: In robotics, Dual Quaternions ($\mathbb{DH}$) are often used to model rigid body motions (screws). It is important to note that Dual Quaternions are also a subalgebra of CGA. Specifically, the even subalgebra of $\mathcal{C}l_{4,1}$ contains elements isomorphic to dual quaternions. However, existing “Dual Quaternion RNNs” often implement the algebra via $4\times 4$ matrix representations or dual-number libraries that introduce significant computational overhead. Versor’s bit-masked formulation offers a more efficient execution model for this same algebraic structure.